OpenStudy (anonymous):

write the slope of the line? (5,0)(5,8)(5,-8)

OpenStudy (anonymous):

\[slope = \frac{y_2-y_1}{x_2-x_1} = \frac{8-0}{5-5} = \frac{8}{0} = undefined\] Basically the line is going straight up and down, so its slope is undefined.